OpenStudy (anonymous):

I need an A in Algebra! Somebody please help!! [10.07] Trevor is tiling his bathroom floor, which has an area that is represented as 175r5 square inches. Each tile has an area of one hundred seventy five r to the fifth power, all over the square root of the quantity sixty four r to the seventeenth power. The total number of tiles used can be represented by the expression below. one hundred seventy five r to the fifth power, all over the square root of the quantity sixty four r to the seventeenth power Simplify the expression for the total number of tiles used. Show your work.

OpenStudy (anonymous):

N = (total area)/(area of one tile) = 175r^5/√(64r^17) Simplify the denominator: √(64r^17) = √(64) √(r^17) = 8 √(r^16 * r^1) = 8 √(r^16) √r = 8 r^8 √r Putting this into the equation for N, N = 175r^5 / (8 r^8 √r) = (175/8) / (r^3 √r) = (175/8) / (r^3 r^½) = (175/8) /r^(7/2) which may also be written = (175/8) r^(-7/2)